OpenStudy (anonymous):

How long will it take $1800 to triple if it is invested at 10% interest compounded continuously? (Express your answer to the nearest tenth.)

OpenStudy (anonymous):

5400 = 1800(1.10)^t Isolate t. 3 = (1.10)^t log 3 = t log 1.10 t = log 3 / log1.10 =11.5years Do the rounding/